OpenStudy (gazedatwrong):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (anonymous):

y = ax + b -3 = 7a + b b = -7a - 3 and -8 = 4a + b b = -4a - 8 then -7a - 3 = -4a - 8 3a = 5 a = 5/3 b = - 4(5/3) - 8 b = - 20/3 - 24/3 = -44/3 y = (5/3)x - 44/3 3y = 5x - 44 5x - 3y = 44 ..

OpenStudy (brooke2353):

For this one you want to do the same thing. You want to find the change in y and x. The change in y is -8-(-3)=-5 And the change in x is 4-7=-3 So your slope is 5/3. Using point slope again you have: y-(-3)=(5/3)(x-7) y+3=5x3−353 Subtracting 3 gives you: y=5x3−443 Putting the x on the left side and multiplying by 3 to get: 3y−5x=−44